# A Cosmic Miracle: A Remarkably Luminous Galaxy at Confirmed with JWST

### v1

@@ 3.1.2 Lyman- break
We perform an independent verification of the emission-line redshift by modeling the observed break as a Ly break at a redshift of . The sharpness of the break leaves little doubt that this is indeed a Ly break â see Fig. 13 in the Appendix for a comparison against a Balmer break spectrum that provides a poor match to the shape (a smooth, gradual decline vs. the sharp edge) and depth of the break (flux would be detected blueward of the break at the depth of these observations).
Here we describe our model set-up to fit the Ly break redshift. The prescription from Miralda-EscudĂ© (1998) is used to approximate the absorption profile due to the Gunn-Peterson effect (Gunn & Peterson, 1965) from neutral hydrogen (Hâi) in the intergalactic medium (IGM). The correction to this profile from Totani et al. (2006) is used to determine the average neutral Hâi fraction, , of the IGM in the line of sight up to the Ly redshift, . We assume an underlying intrinsic galaxy continuum described by a power-law, , where is the rest-frame UV spectral slope and is an arbitrary normalization factor. We use dynesty (Speagle, 2019) to estimate the posteriors on the free parameters in the model (, , , and ) and the total evidence of the distribution. We convolve the model to match the wavelength-dependent spectral resolution of the data.
We find a best-fit Ly break redshift of , consistent with the photometric redshift and the inferred emission-line redshift (see Fig. 2). Intriguingly, the best-fit IGM neutral fraction is at an epoch where virtually all reionization models predict (e.g., Finkelstein et al., 2019; Mason et al., 2019; Naidu et al., 2020b; Hutter et al., 2021; Matthee et al., 2022; Kannan et al., 2022). We also do not require a strong DLA to model this break, which is uncommon at these redshifts (e.g., Curtis-Lake et al., 2023; Carniani et al., 2024; Heintz et al., 2024a). We will explore this in more detail in § 3.3.
@@ 3.2.1 Morphology
The morphology of MoM-z14 is of key interest to understand the origins of its UV luminosity. Among the handful of spectroscopically confirmed bright objects, there appear to be two types â point-sources indistinguishable from the PSF (e.g., GNz11 and GLASS-z12/GHz2), and extended galaxies (e.g., JADES-GS-z14-0 and âMaisieâs Galaxyâ; Finkelstein et al. 2022b) with sizes following empirical scaling relations from lower redshifts (e.g., ; Shibuya et al. 2015). Furthermore, as noticed by Harikane et al. (2024b), these morphological differences are reflected in chemical abundance patterns, signaling a deeper connection between morphology and evolutionary pathways (Fig. 5, and further discussed in §4.3).
@@
We present the forcepho model for MoM-z14 in Fig. 4. The thumbnails shown are 30 mas/pixel images corresponding to 98 pc at . We find the source is resolved and compact (circularized pc; semi-major axis, pc) with a Sersic index of . The source is extended along the North-South direction in the thumbnails shown, with an elongated axis ratio (). There may be a hint of a second component in the images responsible for this elongation. We defer detailed investigation of a two-component model with deeper imaging to future work. We have verified that we find similar results for the size ( pc) by modeling the source with pysersic (Pasha & Miller, 2023) directly on the mosaicked images using empirical PSFs constructed following Weibel et al. (2024).
@@ 3.2.2 SED Modeling
We model the SED using the prospector Bayesian modeling framework (Leja et al., 2017, 2019; Johnson et al., 2021). Our setup closely follows the choices validated in Tacchella et al. (2022); Naidu et al. (2022c, 2024). We use FSPS (Conroy et al., 2009, 2010; Conroy & Gunn, 2010b) with the BPASS stellar models (Stanway & Eldridge, 2018), in particular the v2.2 -bin-imf135all 100 models that assume a Salpeter (1955) IMF with a cutoff. Nebular emission is modeled with the CLOUDY (Ferland et al., 2017) grid produced in Byler et al. (2017). The parameters we fit include seven bins describing a non-parametric star-formation history, the total stellar mass, stellar and gas-phase metallicities, nebular emission parameters, and a flexible dust model (Kriek & Conroy, 2013). We adopt a âbursty continuityâ prior for the star-formation history (Tacchella et al., 2022) with the time bins logarithmically spaced up to a formation redshift of . We hold the first two bins fixed at lookback times of 0-5 and 5-10 Myr to capture bursts that power strong emission lines (e.g., Whitler et al., 2022; Tacchella et al., 2023b) that are increasingly ubiquitous towards the Epoch of Reionization and beyond (e.g., Matthee et al., 2023; Meyer et al., 2024; Covelo-Paz et al., 2025; Lin et al., 2025). Posteriors are sampled using dynesty (Speagle, 2018). It is important to note that systematic uncertainties loom over our model assumptions. For example, it is unclear whether our adopted IMF is applicable to a luminous galaxy (e.g., Cameron et al., 2024; Hutter et al., 2025; Yung et al., 2024). The SED fitting results must therefore be viewed as a baseline set of quantities derived under canonical assumptions.
@@
We fit the model at fixed redshift to the NIRCam photometry and the integrated Câiii] emission line flux. Of all the lines observed, Câiii] is modeled best by the Byler et al. (2017) grid. The other lines are interpreted in the following section using a more flexible model (Cue; Li et al. 2024a) that is capable of capturing, e.g., super-solar [N/O] to produce strong nitrogen lines. The posteriors of the SED and star-formation history (SFH) are shown in Fig. 6. The quality of the fit is reasonable (), with the key area for improvement being the Câiii] line flux that is underestimated by 0.3 dex. We note that this disagreement is worse when using the MIST (Choi et al., 2017, 2020) stellar library with a Chabrier (2003) IMF.
@@
We infer MoM-z14 to be a relatively low-mass galaxy, comparable to the present-day Small Magellanic Cloud (; van der Marel et al. 2009). There is negligible dust attenuation, as evidenced by the blue UV slope (). We are observing this source during a burst phase, wherein the star-formation rate may have increased by up to an order of magnitude in a short span of Myr. We have verified that the high Câiii] EW is driving this rise in the SFH â a fit without this line returns a prior-dominated, relatively flat SFH. Indeed, dedicated grids built to explore Câiii] show that Ă Â EWs signal Myr bursts, high ionization parameters (), and low gas-phase metallicities () (e.g., Jaskot & Ravindranath, 2016; Nakajima et al., 2018). While the galaxy is in the throes of a burst, we cannot rule out the presence of even older stellar populations as reflected in the SFH posteriors, given that we are working purely with the rest-UV. Detailed chemical abundances from high-resolution spectroscopy (e.g., with NIRSpec, ALMA) may be the most efficient path to refining the SFH as the relative contributions of âdelayedâ channels (e.g., neutron star mergers, low-mass stars) and âpromptâ channels (e.g., supermassive stars) manifest in different patterns (e.g., Kobayashi et al., 2020; Johnson et al., 2023; Kobayashi & Ferrara, 2024).
@@ 3.2.3 Emission Line Modeling
We deploy Cue (Li et al., 2024a) to extract insights about the nebular conditions (gas density, ionization parameter), chemical abundances ([O/H], [N/O], [C/O]), and ionizing sources () in MoM-z14. Cue is agnostic to the detailed physics of the source of ionizing photons powering the emission lines and parametrizes the ionizing spectrum as a flexible piecewise power-law. This is particularly useful for modeling luminous objects where the ionizing sources are presently unknown and may not be captured in any ab initio model. Emission lines are predicted via neural net emulation trained on CLOUDY (Ferland et al., 2017). Posteriors are sampled using dynesty (Speagle, 2019). This approach captures the complex covariance between various parameters as opposed to flattening this high-dimensional space into two-dimensional diagnostics, i.e., using specific sets of lines to bracket the density/temperature/metallicity vs. using all the lines at once to constrain all parameters simultaneously (e.g., Li et al., 2024b).
@@
The results from Cue are summarized in Fig. 7 and listed in Table 1. Note that Nâiv is not emulated in Cue and we do not use it in the fitting. All lines are satisfactorily matched within uncertainties, producing an overall . We caution that the posteriors are wide and permit a broad range of outcomes. This is unsurprising given that we are working with a low-resolution spectrum where closely spaced lines are blended (e.g., the Nâiii] quintuplet, Heâii+Oâiii]), and where individual line fluxes are uncertain (). Nonetheless, these results give us a first order, preliminary portrait of the ISM in MoM-z14.
The fits suggest an efficiently ionizing source â â irradiating the gas with an ionization parameter of . Within errors, this is consistent with star-forming galaxies observed at (e.g., Simmonds et al., 2024), with Very Massive/Super Massive Stars that lie around the stellar population maximum (e.g., Schaerer et al., 2025), as well as metal-poor AGN (e.g., Nakajima & Maiolino, 2022). The inferred gas density is high, higher than JADES-GS-z14-0, which was found to be gas depleted perhaps due to intense feedback (e.g., Schouws et al., 2025; Carniani et al., 2025b). Resolving the density sensitive Nâiii], Nâiv], and Câiii] lines with spectra will help us precisely constrain how high the gas density is.
@@
The derived chemical abundance patterns are contextualized in Fig. 8. We find evidence for super-solar [N/O] at relatively low [O/H] of , while the [C/O] is as expected based on local scaling relations. Though note that the [N/O] from Cue may be underestimated given that we do not include Nâiv]. To derive an independent estimate of [N/C], we follow Villar-MartĂn et al. (2004) who performed a detailed analysis of the nitrogen-emitting Lynx arc (Fosbury et al., 2003), a lensed galaxy at . In particular, we can bypass the highly uncertain Oxygen abundance (which we have constraints on only through a blend) and directly derive N/C (N3++N2+)/(C3++C2+) where these ionic abundances follow from NâIV], NâIII], CâIV, and CâIII]. The free parameter is the electron temperature, for which we explore a wide range () bracketing typical values inferred for comparable sources (e.g., Cameron et al., 2023a; CalabrĂČ et al., 2024; Carniani et al., 2025b). This translates to values of to , which maps to to . That is, pending confirmation with higher resolution, higher SNR spectroscopy, MoM-z14 may be among the most nitrogen enhanced galaxies discovered with JWST yet with a enhancement relative to the solar abundance.
@@
Such nitrogen enrichment has been reported in a wide range of objects with JWST â two of the brightest sources at (GNz11, GLASS-z12/GHz2; e.g., Cameron et al. 2023a; Castellano et al. 2024), some Little Red Dots and broad-line AGN (e.g., Ăbler et al., 2023; Labbe et al., 2024; Isobe et al., 2025), and half a dozen individual star-forming galaxies (e.g., Isobe et al., 2023; Topping et al., 2024; Marques-Chaves et al., 2024a; Schaerer et al., 2024). Prior to JWST, only a handful of such strong N-emitting galaxies had been reported across the entire extragalactic observational literature (Fosbury et al., 2003; PatrĂcio et al., 2016; Mingozzi et al., 2022; Pascale et al., 2023). Like MoM-z14, most of these sources are extremely compact relative to their peers at similar redshift, perhaps suggesting shared origins (e.g., Harikane et al., 2024b; Schaerer et al., 2024).
@@
The key to unraveling this shared origin may lie in Galactic archaeology â elevated [N/C] comparable to these sources is routinely observed in Milky Way globular clusters (GCs) as well as some of the most ancient, metal-poor stars born in the Galaxy (e.g., Belokurov & Kravtsov, 2023, 2024). MoM-z14 and the range of objects discussed above may be the âlive actionâ versions of these dense massive clusters and early epochs of star-formation â we discuss this line of reasoning further in §4.3.
@@ 3.3 An Absent Damping Wing at ?
A puzzling discovery from early observations with JWST was a set of sources with apparent strong UV âturnoversâ where the Ly break appeared offset from the systemic redshift. This was first interpreted as extremely strong damped Ly absorption (DLA) due to dense Hâi gas in the vicinity of these galaxies by Heintz et al. (2024a). This phenomenon is now frequently observed among the most distant galaxies at (Umeda et al., 2024; Hainline et al., 2024; DâEugenio et al., 2024; Asada et al., 2024; Witstok et al., 2025; Heintz et al., 2024b, 2025). Contributions from strong nebular continuum and 2-photon emission have also been argued to explain at least a subset of galaxies with strong UV turnovers (Cameron et al., 2024; Katz et al., 2024). However, this interpretation has been excluded for the vast majority of high- galaxies based on the lack of strong emission-line EWs or too blue rest-UV slopes (Chen et al., 2024; Witstok et al., 2025, Pollock et al. in prep.).
@@
In the top panel of Fig. 9 we show qualitatively that MoM-z14 has a relatively sharp Ly break and does not show a broad UV turnover. We contrast the source with JADES-GS-z14-0 at a slightly lower redshift that displays a DLA profile corresponding to (Heintz et al., 2025; Carniani et al., 2024). Not only is such a DLA apparently absent in MoM-z14, but even the IGM around it may be partially ionized. The bottom panel of Figure 9 shows Ly damping wing curves for the IGM with neutral fractions fixed to and 1.0 (i.e. 1% to 100% neutral). Intriguingly, this comparison hints that MoM-z14 may reside in a partially ionized IGM (see also the recent discovery of Ly at in Bunker et al. 2023 and Witstok et al. 2025). We also note that in contrast to JADES-GS-z14-0, MoM-z14âs blue UV slope and inferred hard ionizing spectrum are telltale signs of a non-zero Lyman continuum escape fraction, consistent with it contributing to the ionization of its surroundings, helping explain the difference in the damping signatures between these two sources (e.g., Chisholm et al., 2022; Naidu et al., 2022b; Jaskot et al., 2024a, b). However, we end this section by emphasizing caution that at the moment we are unable to constrain the strength of Ly or NâV emission, which at this low resolution and SNR makes detailed interpretation of the break challenging. We defer fitting IGM+DLA models to future work.

### v2

@@ 3.1.2 Lyman- break
We perform an independent verification of the emission-line redshift by modeling the observed break as a Ly break at a redshift of . The sharpness of the break leaves little doubt that this is indeed a Ly break – see Fig. 14 in the Appendix for a comparison against a typical Balmer break spectrum that provides a poor match to the shape (a smooth, gradual decline vs. the sharp edge) and depth of the break (flux would be detected blueward of the break at the depth of these observations). A recently discovered class of objects – so called “Black Hole Stars” (e.g., Naidu et al., 2025; de Graaff et al., 2025) thought to power the Little Red Dots (e.g., Matthee et al., 2024) – show the largest Balmer breaks on record that may mimic Lyman breaks in broadband photometry, but these sources are excluded based on their characteristic red optical continuum and broad Balmer lines that are not seen in the spectrum.
Here we describe our model set-up to fit the Ly break redshift. The prescription from Miralda-Escudé (1998) is used to approximate the absorption profile due to the Gunn-Peterson effect (Gunn and Peterson, 1965) from neutral hydrogen (H i) in the intergalactic medium (IGM). The correction to this profile from Totani et al. (2006) is used to determine the average neutral H i fraction, , of the IGM in the line of sight up to the Ly redshift, . We assume an underlying intrinsic galaxy continuum described by a power-law, , where is the rest-frame UV spectral slope and is an arbitrary normalization factor. We use dynesty (Speagle, 2019) to estimate the posteriors on the free parameters in the model (, , , and ) and the total evidence of the distribution. We convolve the model to match the wavelength-dependent spectral resolution of the data. Emission lines are fit simultaneously along with the continuum slope.
We find a best-fit Ly break redshift of , consistent with the photometric redshift and the inferred emission-line redshift (see Fig. 2). Intriguingly, the best-fit IGM neutral fraction is non-unity at an epoch where virtually all reionization models predict (e.g., Finkelstein et al., 2019; Mason et al., 2019; Naidu et al., 2020b; Hutter et al., 2021; Matthee et al., 2022; Kannan et al., 2022). We also do not require a strong DLA to model this break, which is uncommon at these redshifts (e.g., Curtis-Lake et al., 2023; Carniani et al., 2024; Heintz et al., 2024b). We will discuss these constraints in detail in § 3.3.
@@ 3.2.1 Morphology
The morphology of MoM-z14 is of key interest to understand the origins of its UV luminosity. Among the handful of spectroscopically confirmed bright objects, there appear to be two types – point-sources indistinguishable from the PSF (e.g., GNz11 and GLASS-z12/GHz2), and extended galaxies (e.g., JADES-GS-z14-0 and “Maisie’s Galaxy”; Finkelstein et al. 2022a) with sizes following empirical scaling relations from lower redshifts (e.g., ; Shibuya et al. 2015). Furthermore, as noticed by Harikane et al. (2024a), these morphological differences are reflected in chemical abundance patterns, signaling a deeper connection between morphology and evolutionary pathways (Fig. 5, and further discussed in §4.3).
@@
We present the forcepho model for MoM-z14 in Fig. 4. The thumbnails shown are 30 mas/pixel images corresponding to 98 pc at . We find the source is resolved and compact (circularized pc; semi-major axis, pc) with a Sersic index of . The source is extended along the North-South direction in the thumbnails shown, with an elongated axis ratio (). Further evidence for this elongated morphology is seen in the radial surface brightness profiles that are extended relative to the PSF (see Fig. 16 in Appendix). There may be a hint of a second component in the images responsible for this elongation. We defer detailed investigation of a two-component model with deeper imaging to future work. We have verified that we find similar results for the size ( pc) by modeling the source with pysersic (Pasha and Miller, 2023) directly on the mosaicked images using empirical PSFs constructed following Weibel et al. (2024).
@@ 3.2.2 SED Modeling
We model the SED using the prospector Bayesian modeling framework (Leja et al., 2017, 2019; Johnson et al., 2021). Our setup closely follows the choices validated in Tacchella et al. (2022); Naidu et al. (2022c, 2024). We use FSPS (Conroy et al., 2009, 2010; Conroy and Gunn, 2010b) with the BPASS stellar models (Stanway and Eldridge, 2018), in particular the v2.2 -bin-imf135all 100 models that assume a Salpeter (1955) IMF with a cutoff. Nebular emission is modeled with the CLOUDY (Ferland et al., 2017) grid produced in Byler et al. (2017). The parameters we fit include seven bins describing a non-parametric star-formation history, the total stellar mass, stellar and gas-phase metallicities, nebular emission parameters, and a flexible dust model (Kriek and Conroy, 2013). We adopt a “bursty continuity” prior for the star-formation history (Tacchella et al., 2022) with the time bins logarithmically spaced up to a formation redshift of . We hold the first two bins fixed at lookback times of 0-5 and 5-10 Myr to capture bursts that power strong emission lines (e.g., Whitler et al., 2022; Tacchella et al., 2023b) that are increasingly ubiquitous towards the Epoch of Reionization and beyond (e.g., Matthee et al., 2023; Meyer et al., 2024; Covelo-Paz et al., 2025; Lin et al., 2025). Posteriors are sampled using dynesty (Speagle, 2018). It is important to note that systematic uncertainties loom over our model assumptions. For example, it is unclear whether our adopted IMF is applicable to a luminous galaxy (e.g., Cameron et al., 2024; Hutter et al., 2025; Yung et al., 2024). The SED fitting results must therefore be viewed as a baseline set of quantities derived under canonical assumptions.
@@
We fit the model at fixed redshift to the NIRCam photometry and the integrated C iii] emission line flux. Of all the lines observed, C iii] is modeled best by the Byler et al. (2017) grid. The other lines are interpreted in the following section using a more flexible model (Cue; Li et al. 2024a) that is capable of capturing, e.g., super-solar [N/O] to produce strong nitrogen lines. The posteriors of the SED and star-formation history (SFH) are shown in Fig. 6. The quality of the fit is reasonable (), with the key area for improvement being the C iii] line flux that is underestimated by 0.3 dex. We note that this disagreement is worse when using the MIST (Choi et al., 2017, 2020) stellar library with a Chabrier (2003) IMF.
@@
We infer MoM-z14 to be a relatively low-mass galaxy, comparable to the present-day Small Magellanic Cloud (; van der Marel et al. 2009). There is negligible dust attenuation, as evidenced by the blue UV slope (). We are observing this source during a burst phase, wherein the star-formation rate may have increased by up to an order of magnitude in a short span of Myr. We have verified that the high C iii] EW is driving this rise in the SFH – a fit without this line returns a prior-dominated, relatively flat SFH. Indeed, dedicated grids built to explore C iii] show that Å EWs signal Myr bursts, high ionization parameters (), and low gas-phase metallicities () (e.g., Jaskot and Ravindranath, 2016; Nakajima et al., 2018). While the galaxy is in the throes of a burst, we cannot rule out the presence of even older stellar populations as reflected in the SFH posteriors, given that we are working purely with the rest-UV. Detailed chemical abundances from high-resolution spectroscopy (e.g., with NIRSpec, ALMA) may be the most efficient path to refining the SFH as the relative contributions of “delayed” channels (e.g., neutron star mergers, low-mass stars) and “prompt” channels (e.g., supermassive stars) manifest in different patterns (e.g., Kobayashi et al., 2020; Johnson et al., 2023; Kobayashi and Ferrara, 2024).
@@ 3.2.3 Emission Line Modeling
We deploy Cue (Li et al., 2024a) to extract insights about the nebular conditions (gas density, ionization parameter), chemical abundances ([O/H], [N/O], [C/O]), and ionizing sources () in MoM-z14. Cue is agnostic to the detailed physics of the source of ionizing photons powering the emission lines and parametrizes the ionizing spectrum as a flexible piecewise power-law. This is particularly useful for modeling luminous objects where the ionizing sources are presently unknown and may not be captured in any ab initio model. Emission lines are predicted via neural net emulation trained on CLOUDY (Ferland et al., 2017). Note that Cue is designed to model emission lines, and not to perform full spectrum fitting. Posteriors are sampled using dynesty (Speagle, 2019). This approach captures the complex covariance between various parameters as opposed to flattening this high-dimensional space into two-dimensional diagnostics, i.e., using specific sets of lines to bracket the density/temperature/metallicity vs. using all the lines at once to constrain all parameters simultaneously (e.g., Li et al., 2024b).
@@
The results from Cue are summarized in Fig. 7 and listed in Table 1. In Fig. 7 we have summarized the ionizing spectrum in the form of the familiar for simplicity – in Fig. 15 we show all the individual power law parameters. Note that N iv is not emulated in Cue and we do not use it in the fitting. All lines are satisfactorily matched within uncertainties, producing an overall . We caution that the posteriors are wide and permit a broad range of outcomes. This is unsurprising given that we are working with a low-resolution spectrum where closely spaced lines are blended (e.g., the N iii] quintuplet, He ii+O iii]), and where individual line fluxes are uncertain (). Further, posteriors for almost all quantities are truncated at the edges of the grid that Cue is trained on. This is an extensive grid with wider coverage than typical photoionization models (e.g., Feltre et al., 2016; Byler et al., 2017), and yet extreme sources like MoM-z14 may require even broader sets of parameters. Nonetheless, the results in Fig. 7 give us a first order, preliminary portrait of the ISM in MoM-z14.
The Cue fits that take into account the observed EWs and line ratios suggest an efficiently ionizing source – – irradiating the gas with an ionization parameter of . Within errors, this is consistent with star-forming galaxies observed at (e.g., Simmonds et al., 2024), with Very Massive/Super Massive Stars that lie around the stellar population maximum (e.g., Schaerer et al., 2025), as well as metal-poor AGN (e.g., Nakajima and Maiolino, 2022). The relatively uncertain measurements currently in hand are unable to directly isolate an ionizing source across these scenarios – the He 2, C 3], C 4, N 4], N 3] EWs are all consistent with metal-poor AGN as well as star-forming galaxies. Indeed, Calabrò et al. (2024) have shown in detail the difficulty of distinguishing between these populations at PRISM resolution using EWs of these emission lines in their analysis of GLASS-z12/GHz2. However, as we discuss later here and in §4.3, the chemical abundances, especially [N/C] are relatively robust among all our inferred quantities and may already yield insights on the ionizing sources.
The inferred gas density is higher than JADES-GS-z14-0, which was found to be gas depleted perhaps due to intense feedback (e.g., Schouws et al., 2025; Carniani et al., 2025b). Resolving the density sensitive N iii], N iv], and C iii] lines with spectra will help us precisely constrain exactly how high the gas density is. For now we note that while high ionization parameters () and electron densities (¿ cm-3) might seem extreme, such values are plausible for high-redshift galaxies (e.g., Topping et al., 2024) and may even indicate AGN activity (e.g., Maiolino et al., 2024).
@@
The derived chemical abundance patterns are contextualized in Fig. 8. We find evidence for super-solar [N/O] at relatively low [O/H] of , while the [C/O] is as expected based on local scaling relations. Though note that the [N/O] from Cue may be underestimated given that we do not include N iv]. To derive an independent estimate of [N/C], we follow Villar-Martín et al. (2004) who performed a detailed analysis of the nitrogen-emitting Lynx arc (Fosbury et al., 2003), a lensed galaxy at . In particular, we can bypass the highly uncertain Oxygen abundance (which we have constraints on only through a blend) and directly derive N/C (N3++N2+)/(C3++C2+) where these ionic abundances follow from N IV], N III], C IV, and C III]. The free parameter is the electron temperature, for which we explore a wide range () bracketing typical values inferred for comparable sources (e.g., Cameron et al., 2023a; Calabrò et al., 2024; Carniani et al., 2025b). This translates to values of to , which maps to to . That is, pending confirmation with higher resolution, higher SNR spectroscopy, MoM-z14 may be among the most nitrogen enhanced galaxies discovered with JWST yet with a enhancement relative to the solar abundance.
@@
Such nitrogen enrichment has been reported in a wide range of objects with JWST – two of the brightest sources at (GNz11, GLASS-z12/GHz2; e.g., Cameron et al. 2023a; Castellano et al. 2024), some Little Red Dots and broad-line AGN (e.g., Übler et al., 2023; Labbe et al., 2024; Isobe et al., 2025), and half a dozen individual star-forming galaxies (e.g., Isobe et al., 2023; Topping et al., 2024; Marques-Chaves et al., 2024; Schaerer et al., 2024). Prior to JWST, only a handful of such strong N-emitting galaxies had been reported across the entire extragalactic observational literature (Fosbury et al., 2003; Patrício et al., 2016; Mingozzi et al., 2022; Pascale et al., 2023). Like MoM-z14, most of these sources are extremely compact relative to their peers at similar redshift, perhaps suggesting shared origins (e.g., Harikane et al., 2024a; Schaerer et al., 2024).
@@
The key to unraveling this shared origin may lie in Galactic archaeology – elevated [N/C] comparable to these sources is routinely observed in Milky Way globular clusters (GCs) as well as some of the most ancient, metal-poor stars born in the Galaxy (e.g., Belokurov and Kravtsov, 2023, 2024). MoM-z14 and the range of objects discussed above may be the “live action” versions of these dense massive clusters and early epochs of star-formation – we discuss this line of reasoning further in §4.3.
@@ 3.3 An Absent Damping Wing at ?
A puzzling discovery from early observations with JWST was a set of sources with apparent strong UV “turnovers” where the Ly break appeared offset from the systemic redshift. This was first interpreted as extremely strong damped Ly absorption (DLA) due to dense H i gas in the vicinity of these galaxies by Heintz et al. (2024b). This phenomenon is now frequently observed among the most distant galaxies at (Umeda et al., 2024; Hainline et al., 2024; D’Eugenio et al., 2024; Asada et al., 2024; Witstok et al., 2025; Heintz et al., 2024a, 2025). Contributions from strong nebular continuum and 2-photon emission have also been argued to explain at least a subset of galaxies with strong UV turnovers (Cameron et al., 2024; Katz et al., 2024). However, this interpretation has been excluded for the vast majority of high- galaxies based on the lack of strong emission-line EWs or too blue rest-UV slopes (Chen et al., 2024; Witstok et al., 2025, Pollock et al. in prep.).
@@
In the top panel of Fig. 9 we show qualitatively that MoM-z14 has a relatively sharp Ly break and does not show a broad UV turnover. We contrast the source with JADES-GS-z14-0 at a slightly lower redshift that displays a DLA profile corresponding to (Heintz et al., 2025; Carniani et al., 2024). Not only is such a DLA apparently absent in MoM-z14, but even the IGM around it may be partially ionized. The bottom panel of Figure 9 shows Ly damping wing curves for the IGM with neutral fractions fixed to and 1.0 (i.e. 1% to 100% neutral). Intriguingly, this comparison hints that MoM-z14 may reside in a partially ionized IGM (see also the recent discovery of Ly at in Bunker et al. 2023 and Witstok et al. 2025). We also note that in contrast to JADES-GS-z14-0, MoM-z14’s blue UV slope and inferred hard ionizing spectrum are telltale signs of a non-zero Lyman continuum escape fraction, consistent with it contributing to the ionization of its surroundings, helping explain the difference in the damping signatures between these two sources (e.g., Chisholm et al., 2022; Naidu et al., 2022a; Jaskot et al., 2024a, b).
To quantify these insights, we fit a model similar to the one discussed in §3.1.2 but with the redshift fixed to the systemic UV line redshift () and an extra parameter () so we may account for the local, dense neutral gas (i.e., the DLA) in addition to the neutral fraction of the IGM (). The DLA adds an additional Ly absorption component to the power-law+IGM model introduced in Sect. 3.1.2 in the form of a Voigt profile, using the approximation from Tepper-García (2006).
In Fig. 10 we present the results of the fit to this model. The posteriors are broad, but bear intriguing hints of an early ionized region. A strong DLA as seen in JADES-GS-z14-0 () and typical galaxies (; Heintz et al. 2024a) is disfavored at confidence. The IGM neutral fraction is inferred to be with a completely neutral sight-line disfavored at confidence. We emphasize caution that at the moment we are unable to constrain the strength of Ly or N V emission, which at this low resolution and SNR makes detailed interpretation of the break challenging. However, we hope these preliminary results motivate the community to undertake the intensive spectroscopy required to definitively collapse these posteriors.
It is interesting to note that even a modest ionized fraction at may help alleviate some of the most pressing tensions in CDM cosmology (e.g., the problem of negative neutrino masses and hints of evolving dark energy) by providing support for a higher value of the Thomson optical depth () compared to the value inferred by Planck Collaboration et al. (2018) (e.g., Giarè et al., 2024; Asthana et al., 2024; Mason et al., 2025; Sailer et al., 2025).